OpenStudy (anonymous):

Algebra 2 3^(x-5)=9 Could anyone please explain to me how to solve this? a) x=-3 b)x=2 c)x=7 d)x=8

jimthompson5910 (jim_thompson5910):

3^(x-5)=9 3^(x-5)=3^2 The bases are both equal to 3, so the exponents must be equal, which means x-5 = 2 Solve for x to get your answer

OpenStudy (anonymous):

Oh! I forgot to break down 9 and I thought it was x-5=1 and that wasn't one of the choices. Thank you!

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

That's a nice little trick if you're able to notice something like that. The other thing you can do here is take log base 3 of both sides. You can think of log base 3 as being the opposite of 3^() so the operations will cancel out. On the left side: \[\Large \log_{3} (3^{x+5}) = x+5\] On the right side: \[\Large \log_{3}9 = 2 \] And so you get x+5 = 2

OpenStudy (anonymous):

Thank you! I will keep that in mind if that ever shows up again. (: